# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 8 Conclusions
This paper presented the results of a SLR on the application of spatio-temporal GNN models to time series classification and forecasting problems in different fields. Lately, GNNs have gained significant popularity due to their ability to process graph-structured data. This has led in more recent years to the development of spatio-temporal GNNs in the field of time series analysis, due to their ability to model dependencies between variables and across time points.
This SLR has brought forward on two set of questions: generic questions which can be answered at the level of bibliographic overview, and more specific questions regarding particular aspects of the proposed spatio-temporal GNN models. By answering these questions, it has emerged that the majority of the models in the selected papers is characterized by a convolutional spatial aggregation. However, several graph attentional models are also emerging. In addition, while the majority of the selected papers focus on models with a pre-defined graph structure, an increasing number of studies, particularly in the ”Generic” group, are beginning to develop models that learn the graph structure autonomously.
A first key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig. 4. This can be attributed to the fact that the involved researchers come from different communities with focus on specific application domains. As a result, there seems to be a lack of standardised datasets or benchmarks. The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide a foundation for future studies. A second major key point resulting from the proposed SLR seems hence to be that there is a need for the GNN research community to work on common datasets and develop comparable and reproducible models. This would enhance transparency and make it easier to evaluate advancements in the field and to share them more easily.
@@ Appendix A List of selected journal papers
Tab. LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
Table 35: List of selected journal papers, with year, group they belong to, case study, and nature of the task.
@@
[art_25] 2022 Energy Urban buildings energy consumption Forecasting
@@
[art_38] 2022 Environment Frost Classification/forecasting
@@
[art_71] 2023 Generic Clinical risk Classification
@@
[art_76] 2023 Generic Long-term series forecasting Forecasting
@@
[art_82] 2023 Generic Multivariate long sequence time series Forecasting
@@
[art_84] 2022 Generic Multivariate time series Classification/forecasting
@@
[art_88] 2023 Generic Multi-channel time series Forecasting
@@
[art_202] 2021 Generic Generic Classification/forecasting
@@
[art_95] 2023 Mobility Flight delays, urban traffic Forecasting
@@
[art_104] 2022 Mobility Internet of vehicles Forecasting
@@
[art_12] 2023 Other Modulation classification Classification
[art_13] 2022 Other Modulation classification Classification
[art_118] 2023 Other Encrypted traffic classification Classification
[art_136] 2022 Other Spatial temporal events Forecasting
[art_137] 2023 Other Quality indicators Forecasting
[art_138] 2023 Other Cellular traffic Forecasting
[art_139] 2023 Other Seismic data Forecasting
[art_140] 2023 Other Structural dynamics on irregular domains Forecasting
[art_141] 2023 Other Tailing dam monitoring Forecasting
[art_142] 2023 Other Video engagement Forecasting
[art_143] 2023 Other Subsurface production Forecasting
[art_144] 2023 Other Cellular traffic Forecasting
[art_145] 2023 Other Data centers maintenance Forecasting
[art_146] 2023 Other Hydraulic runoff Forecasting
[art_147] 2022 Other App popularity Forecasting
[art_148] 2021 Other Cellular traffic Forecasting
[art_149] 2022 Other Water demand Forecasting
[art_151] 2023 Other User preference Edge prediction
[art_159] 2021 Other Human activity recognition Classification
[4] 2023 Other Modulation classification Classification

### v2

@@ 8 Conclusions
This paper presented the results of a SLR on the application of spatio-temporal GNN models to time series classification and forecasting problems in different fields. The interest on this subject comes from the fact that lately, GNNs have gained significant popularity due to their ability to process graph-structured data. This has led in more recent years to the development of spatio-temporal GNNs and their application to time series analysis, due to their ability to model dependencies between variables and across time points at once.
This SLR has brought forward two sets of research questions: generic questions which can be answered at the level of a bibliographic overview, and more specific questions regarding particular aspects of the proposed spatio-temporal GNN models. The answers to these questions provide a list of interesting considerations.
It has emerged that the majority of the models in the selected papers is characterized by a convolutional spatial aggregation. However, graph attentional models are also emerging as important tools. In addition, while the majority of the selected papers focus on models with a pre-defined graph structure, an increasing number of studies, particularly in the “Generic” thematic group, are beginning to develop models that learn the graph structure autonomously.
Another key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig. 3. This can be attributed to the fact that the involved researchers come from different communities which focus on distinct application domains. As a result, there seems to be a lack of standardized datasets or benchmarks in the studies. The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide common background information for future studies.
A further major key point resulting from our review, which we believe should be emphasized in these conclusions, is the need for the spatio-temporal GNN research community to work on common datasets and develop comparable and reproducible models. That would enhance transparency, and it would make it easier for all researchers to reciprocally evaluate advancements in the field and to share them more easily.
Supplementary material
Link to the GitHub repository with the links to the 281 reviewed papers, organized by groups: https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git.
CRediT author statement
Flavio Corradini: Project administration, Funding acquisition. Flavio Gerosa: Investigation, Writing - Review & Editing, Visualization. Marco Gori: Writing - Review & Editing, Supervision. Carlo Lucheroni: Writing - Review & Editing, Supervision. Marco Piangerelli: Conceptualization, Methodology, Writing - Review & Editing, Supervision. Martina Zannotti: Conceptualization, Methodology, Validation, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization.
Acknowledgments
This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 2, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041 - VITALITY - CUPJ13C22000430001; and the NextGenerationEU - National Recovery and Resilience Plan (PNRR) grant CUP J11J23001450006, with the contribution of Syeew srl.
@@ Appendix A List of selected journal papers
Tab. LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (classification or forecasting).
Table 41: List of the 281 selected journal papers, with year, group they belong to, case study, and nature of the task.
@@
[art_25] 2022 Energy Urban energy consumption Forecasting
@@
[art_332] 2024 Energy Wind power Forecasting
[art_333] 2024 Energy Wind power Forecasting
[art_334] 2024 Energy Air conditioning control Forecasting
[art_335] 2024 Energy Air conditioning control Forecasting
[art_337] 2024 Energy Electrical demand and prices Forecasting
[art_338] 2024 Energy Electric load Forecasting
[art_339] 2024 Energy Electric load Forecasting
[art_340] 2024 Energy Chiller energy consumption Forecasting
[art_341] 2024 Energy Electricity demand Forecasting
[art_342] 2024 Energy Electric power anomaly detection Classification
[art_343] 2024 Energy Electricity demand Forecasting
[art_344] 2024 Energy Electricity consumption Forecasting
[art_345] 2024 Energy District heating networks Forecasting
[art_347] 2024 Energy Power quality disturbances Classification
[art_348] 2024 Energy Residential load Forecasting
[art_349] 2024 Energy Photovoltaic energy Forecasting
[art_350] 2024 Energy Electric vehicle charging stations Forecasting
[art_351] 2024 Energy Natural gas consumption Forecasting
[art_353] 2024 Energy Networked microgrids Forecasting
[art_390] 2024 Energy Multivariate time series Forecasting
@@
[art_38] 2022 Environment Frost Both
@@
[art_354] 2024 Environment PM2.5 prediction Forecasting
[art_355] 2024 Environment Sea clutter Forecasting
[art_356] 2024 Environment Earthquakes Forecasting
[art_357] 2024 Environment Plankton dynamics Forecasting
[art_359] 2024 Environment Groundwater data Forecasting
[art_360] 2024 Environment Carbon emission Forecasting
[art_361] 2024 Environment PM2.5 prediction Forecasting
[art_363] 2024 Environment Water demand Forecasting
[art_364] 2024 Environment Water quality and flow rate Forecasting
[art_365] 2024 Environment Weather Forecasting
[art_366] 2024 Environment Wind speed scenario Forecasting
[art_367] 2024 Environment Carbon emission Forecasting
[art_368] 2024 Environment Air quality Forecasting
[art_370] 2024 Environment Temperature Forecasting
[art_371] 2024 Environment Air quality Forecasting
[art_373] 2024 Environment Greenhouse gas concentrations Forecasting
@@
[art_374] 2024 Finance Oil and metal market Classification
[art_375] 2024 Finance Supply chain demand Forecasting
[art_377] 2024 Finance Stock price Classification
[art_378] 2024 Finance Stock market volatility Forecasting
[art_379] 2024 Finance Agricultural price Forecasting
[art_381] 2024 Finance Cryptocurrencies’ price Forecasting
[art_382] 2024 Finance Stock trend Classification
@@
[art_71] 2023 Generic Clinical risk classification Classification
@@
[art_76] 2023 Generic Long term series forecasting Forecasting
@@
[art_82] 2023 Generic Multivariate long sequences Forecasting
@@
[art_84] 2022 Generic Multivariate time series Both
@@
[art_88] 2023 Generic Multivariate time series Forecasting
@@
[art_202] 2021 Generic Generic Both
[art_383] 2024 Generic Multivariate time series Forecasting
[art_384] 2024 Generic Multivariate time series Forecasting
[art_385] 2024 Generic Multivariate time series Classification
[art_386] 2024 Generic Multivariate time series Classification
[art_387] 2024 Generic Multivariate time series Both
[art_388] 2024 Generic Multivariate time series Forecasting
[art_389] 2024 Generic Multivariate time series Classification
[art_393] 2024 Generic Multivariate time series Forecasting
[art_395] 2024 Generic Multivariate time series Classification
[art_396] 2024 Generic Multivariate time series Forecasting
[art_397] 2024 Generic Multivariate time series Forecasting
[art_398] 2024 Generic Multivariate time series Forecasting
[art_400] 2024 Generic Multivariate time series Forecasting
[art_401] 2024 Generic Multivariate time series Forecasting
[art_403] 2024 Generic Multivariate time series Both
[art_404] 2024 Generic Multivariate time series Forecasting
[art_405] 2024 Generic Multivariate time series Forecasting
[art_407] 2024 Generic Multivariate time series Classification
[art_417] 2024 Generic Multivariate time series Forecasting
[art_464] 2024 Generic Multivariate time series Forecasting
@@
[art_410] 2024 Health Infectious disease Forecasting
[art_411] 2024 Health EEG emotion recognition Classification
[art_412] 2024 Health EEG emotion recognition Classification
[art_413] 2024 Health Epidemic Forecasting
[art_414] 2024 Health Epidemic Forecasting
[art_415] 2024 Health Epidemic Forecasting
[art_416] 2024 Health Epidemic Forecasting
@@
[art_104] 2022 Mobility Internet of vehicles traffic Forecasting
@@
[art_406] 2024 Mobility Multivariate time series Forecasting
[art_408] 2024 Mobility Multivariate time series Forecasting
[art_419] 2024 Mobility Urban traffic Forecasting
[art_420] 2024 Mobility Urban traffic Forecasting
[art_421] 2024 Mobility Metro flow traffic Forecasting
[art_422] 2024 Mobility Urban traffic Forecasting
[art_424] 2024 Mobility Urban traffic Forecasting
[art_425] 2024 Mobility Network traffic prediction Forecasting
[art_427] 2024 Mobility Urban traffic Forecasting
[art_430] 2024 Mobility Urban traffic Forecasting
[art_431] 2024 Mobility Urban traffic Forecasting
[art_432] 2024 Mobility Urban traffic Forecasting
[art_433] 2024 Mobility Urban traffic Forecasting
[art_434] 2024 Mobility Airport performances Forecasting
[art_435] 2024 Mobility Urban traffic Forecasting
[art_438] 2024 Mobility Urban traffic Forecasting
[art_439] 2024 Mobility Urban traffic Forecasting
[art_441] 2024 Mobility Urban traffic Forecasting
[art_442] 2024 Mobility Urban traffic Forecasting
[art_443] 2024 Mobility Flight delay prediction Forecasting
[art_444] 2024 Mobility Urban traffic Forecasting
[art_445] 2024 Mobility Urban traffic Forecasting
[art_446] 2024 Mobility Bike sharing demand Forecasting
[art_447] 2024 Mobility Urban traffic Forecasting
[art_448] 2024 Mobility Urban traffic Forecasting
[art_449] 2024 Mobility Human mobility Forecasting
[art_451] 2024 Mobility Urban traffic Forecasting
[art_452] 2024 Mobility Bus station Profiling
[art_453] 2024 Mobility Urban traffic Forecasting
[art_454] 2024 Mobility Bus stop mobility Forecasting
[art_12] 2023 Other topics Modulation classification Classification
[art_13] 2022 Other topics Modulation classification Classification
[art_118] 2023 Other topics Encrypted traffic classification Classification
[art_136] 2022 Other topics Spatio-temporal events Forecasting
[art_137] 2023 Other topics Quality indicators prediction Forecasting
[art_138] 2023 Other topics Cellular traffic prediction Forecasting
[art_139] 2023 Other topics Seismic data Forecasting
[art_140] 2023 Other topics Structural dynamics Forecasting
[art_141] 2023 Other topics Tailing dam monitoring Forecasting
[art_142] 2023 Other topics Educational videos engagement Forecasting
[art_143] 2023 Other topics Subsurface production Forecasting
[art_144] 2023 Other topics Cellular Network Traffic Forecasting
[art_145] 2023 Other topics Data centers maintenance Forecasting
[art_146] 2023 Other topics Hydraulic runoff Forecasting
[art_147] 2022 Other topics App popularity Forecasting
[art_148] 2021 Other topics Cellular Network Traffic Forecasting
[art_149] 2022 Other topics Water demand Forecasting
[art_151] 2023 Other topics User preference Edge prediction
[art_159] 2021 Other topics Human activity recognition Classification
[art_303] 2023 Other topics Modulation classification Classification
[art_455] 2024 Other topics Polypropylene production data Forecasting
[4] 2024 Other topics Botnet detection Classification
[art_457] 2024 Other topics Cellular KPI prediction Forecasting
[art_458] 2024 Other topics Tungsten floatation plant Forecasting
[art_459] 2024 Other topics Modulation recognition Classification
[art_460] 2024 Other topics IoT intrusion detection Classification
[art_463] 2024 Other topics Modulation recognition Classification
[art_465] 2024 Other topics Crime prediction Forecasting
[art_468] 2024 Other topics Mobile traffic prediction Forecasting
@@
[art_461] 2024 Predictive monitoring Motor temperature Forecasting
[art_469] 2024 Predictive monitoring Fault diagnosis Classification
[art_470] 2024 Predictive monitoring Root cause localization Forecasting
[art_471] 2024 Predictive monitoring Mechanical fault diagnosis Classification
[art_474] 2024 Predictive monitoring Mechanical fault diagnosis Classification
[art_475] 2024 Predictive monitoring Outlier detection Classification
[art_477] 2024 Predictive monitoring Anomaly detection Forecasting
[art_478] 2024 Predictive monitoring Remaining useful life Forecasting
[art_479] 2024 Predictive monitoring Anomaly detection Classification
[art_480] 2024 Predictive monitoring Sensor data Forecasting
[art_481] 2024 Predictive monitoring Anomaly detection Classification
[art_482] 2024 Predictive monitoring Process factors Forecasting
[art_483] 2024 Predictive monitoring Anomaly detection Classification
[art_484] 2024 Predictive monitoring Anomaly detection Classification
[art_486] 2024 Predictive monitoring Fault diagnosis Classification
[art_487] 2024 Predictive monitoring Fault diagnosis Classification
[art_488] 2024 Predictive monitoring Remaining useful life Forecasting
[art_489] 2024 Predictive monitoring Anomaly detection Classification
[art_491] 2024 Predictive monitoring Fault diagnosis Classification